OpenStudy (anonymous):

A line goes through the points (2, b) and (c, 10). Find the y-coordinate of the point on the line that has an x-coordinate of 3.

OpenStudy (anonymous):

The answer is in terms of c and b.

OpenStudy (accessdenied):

You would find the slope between the two points (we don't know b or c, so we'd just leave them as is), and then insert one of the points along with the slope into point-slope form of an equation. \[ \begin{split} y - y_1 = m(x - x_1)\\ m = \frac{y_2 - y_1}{x_2 - x_1}\\ (x_1,y_1), (x_2, y_2)\\ \end{split} \]

OpenStudy (accessdenied):

* that would define the line itself.. then we'd insert the x=3 to that equation and solve for y

OpenStudy (anonymous):

Will you show me?

OpenStudy (accessdenied):

sure -- you need me to go through the whole problem step by step or a specific part?

OpenStudy (accessdenied):

This problem would play out much like any other "find the equation of a line between these points" and "find y at x=..." problem, except we're using two letters to represent numbers. \[ \large \begin{split} m &= \frac{y_2 - y_1}{x_2 - x_1}\\ &= \frac{10 - b}{c - 2}\\ \end{split} \]

OpenStudy (accessdenied):

\[ \large \begin{split} y - y_1 &= m(x - x_1)\\ y - b &= (\frac{10 - b}{c - 2})(x - 2)\\ y &= (\frac{10 - b}{c - 2})(x - 2) + b\\ \end{split} \] Does that all make sense?

OpenStudy (anonymous):

yes, I got there but wasn't sure what to do next

OpenStudy (accessdenied):

Well, we want to find y when x=3. So, we can just substitute that into our equation and solve (I'd also get one fraction by using common denominators) \[ \large \begin{split} y &= (\frac{10 - b}{c - 2})(3 - 2) + b\\ &= (\frac{10 - b}{c - 2}) + b\\ &= (\frac{10 - b}{c - 2}) + b(\frac{c - 2}{c - 2})\\ &= (\frac{10 - b}{c - 2}) + \frac{bc - 2b}{c - 2}\\ &= (\frac{(10 - b) + (bc - 2b)}{c - 2}\\ &= \frac{10 - 3b + bc}{c - 2}\\ y &= \frac{bc - 3b + 10)}{c - 2}\\ \end{split} \]

OpenStudy (anonymous):

Oh! That's what I didn't do, I didn't multiply b by c-2/c-2. Thanks acc and sorry for being away for awhile.

OpenStudy (accessdenied):

it's alright. as long as you got it now. :)